# Community Forensics: Using Thousands of Generators to Train Fake Image Detectors

### v1

@@ Abstract
One of the key challenges of detecting AI-generated images is spotting images that have been created by previously unseen generative models. We argue that the limited diversity of the training data is a major obstacle to addressing this problem, and we propose a new dataset that is significantly larger and more diverse than prior work. As part of creating this dataset, we systematically download thousands of text-to-image latent diffusion models and sample images from them. We also collect images from dozens of popular open source and commercial models. The resulting dataset contains 2.7M images that have been sampled from 4803 different models. These images collectively capture a wide range of scene content, generator architectures, and image processing settings. Using this dataset, we study the generalization abilities of fake image detectors. Our experiments suggest that detection performance improves as the number of models in the training set increases, even when these models have similar architectures. We also find that detection performance improves as the diversity of the models increases, and that our trained detectors generalize better than those trained on other datasets. [^1] https://jespark.net/projects/2024/community_forensics
@@ 1 Introduction
Figure 1: Performance vs. model diversity. We use images sampled from different numbers of open source latent diffusion models in the Community Forensics dataset to train fake image detectors (shown in Fig. 2(a)). As the number of models increases, so does the detector’s performance, even though these models have similar designs and same total number of images. This improvement is largest for test images from out-of-distribution generative model classes, such as pixel-based diffusion models or GANs. For each data point, we sample 10 random model subsets with 100K training images each and report the mean and standard error values.
@@
We hypothesize that the lack of diversity in training datasets is a major source of these shortcomings. Although today’s datasets often contain millions of fake images, they come from a relatively small number of generators (e.g., fewer than 20 models in any given previous work that we are aware of). As a result, this data fails to capture many sources of variation that one might encounter in the wild. These limitations also make it challenging to accurately benchmark performance, since it is easy for cues that work well on one set of generators to fail on others.
To address these problems, we propose Community Forensics, a dataset that is significantly more diverse and comprehensive than those in prior works (Fig. 2). Our dataset contains images generated by: (a) thousands of systematically downloaded open-source latent diffusion models, (b) hand-selected open source models with various architectures, and (c) state-of-the-art commercial models. We use this dataset to conduct a study of generalization in image forensics.
To acquire large numbers of models, we sample images from thousands of text-to-image diffusion models hosted on a popular model-sharing website, Hugging Face [37]. We exploit the fact that these models use a common programming library [124] and thus can be sampled in a standardized way. A large fraction of them are extensions of Stable Diffusion [102], but collectively capture a variety of common model variations, such as in the architecture, image processing, and image content. We also sample images from many other open source models, including GANs [40], autoregressive models [42], and consistency models [117, 81]. To help study how image content affects classification performance, we provide a corresponding set of real images that are designed to resemble the generated images. For example, we condition the text-to-image models using text obtained by captioning our real images.
@@
We use this dataset to study generalization in the generated image detection problem. Our experiments support the hypothesis that increasing the diversity of generative models used in training is important for generalization. Through experiments, we find:
@@
  • Adding more generative models improves performance. Fig. 1 demonstrates the performance of fake image detection when trained on samples from varying numbers of diffusion models. Notably, the performance improves as more models are added, even across different architectures.
  • Including diverse generative model architectures significantly improves results, since classifiers do not fully generalize between generator architectures. Likewise, the performance gain from including large numbers of images from any particular architecture is relatively marginal.
@@ 2 Related work
A number of datasets have been proposed for specifically detecting “deepfake” images containing manipulated faces [103, 105, 28, 68, 72, 64, 134]. Rather than focusing on face manipulation, we address of creating general-purpose methods that can detect images that have been directly produced by generative models. Wang et al. [126] proposed a widely-used dataset of CNN-generated images, mixing images from GANs [59, 60, 10, 132, 14, 94] with other models [104, 12, 71, 21, 11]. This work showed that forensics models generalize between generative models, providing motivation for training on large datasets of diverse generators. However, their classifier was trained on images from a single GAN and was highly sensitive to data augmentation parameters, and more recent work shows that it does not generalize to newer models [17, 91]. Ojha et al. [91] introduced a dataset of recent diffusion models and found that training a linear classifier on CLIP features [100] extracted from ProGAN-generated images performed well. Cozzolino et al. [18] extends this work by studying the performance of CLIP-based detectors on various generative models and datasets. Epstein et al. [35] simulated detecting fake images in an online way by training a detector up to a certain year and testing it on generators released after that year. Zhu et al. [133] collected 1.4M generated images from 8 different generators. These datasets, however, only consider a handful of models (less than 20 each), limiting the generalization of their detectors. We improve upon these works by collecting much more diverse generative models to improve the performance and generalization of the detector. In concurrent work, Hong et al. [50] acquires user-created images from Midjourney and CivitAI. This strategy is complementary to ours: while it aims to collect in-the-wild fake images, its distribution is centered on images that users share, and the models are not necessarily identifiable, making it challenging to rigorously analyze the dataset’s contents and to interpret experiments conducted on it.
@@
Classic work on image forensics relied on methods based on image statistics [99] and physical constraints [57], rather than learning. A number of datasets have been created for detecting images that have been manipulated using traditional methods, such as with photo editors [29, 54, 24, 65, 88]. Recent works focus on detecting synthetic images by inspecting the generator fingerprints. Zhang et al. [131] and Marra et al. [82] proposed identifying the spatial fingerprints left by the generator to detect synthetic images. Others focus on spectral anomalies to detect synthetic images. Durall et al. [32] and Dzanic et al. [33] identified that CNN-generated images fail to reproduce certain spectral properties of real images. Corvi et al. [17] studies the frequency fingerprints of the generated images and analyzes the cross-architecture generalization of the detector. Bammey [9] uses high-frequency artifacts to detect generated images. However, these approaches may be brittle since the artifacts they rely on can be eliminated by post-processing [18]. We instead approach this problem in a data-driven manner, scaling the number of models, images, and architectures. Recent work has created ensembles of fake image classifiers [51]. In parallel, researchers have detected text generated by language models using supervised learning and heuristics [86, 56, 8, 122, 38, 106, 66, 116], which closely resemble those in visual forensics. However, no existing techniques that we are aware of aim to collect comprehensive datasets of community-created generators.
@@
Our work is related to the out-of-distribution recognition problem as it involves generalizing to unseen generators and image processing. A variety of approaches have been proposed for this problem, based on likelihood ratios [101, 73, 128], self-supervision [112, 87, 48, 125], internal model statistics [47, 107], temperature scaling [6, 74], and via energy-based models [31, 76, 34]. Work by Schuhmann et al. [111] and Hendrycks et al. [49] show that diverse training data and data augmentation is important to improving the robustness to out-of-distribution samples. Our results are in line with these conclusions, as we find that a diverse set of generative models and stronger augmentations improve generalization.
@@ 3 The Community Forensics Dataset
GenImage [133] 8 1.4M Diffusion, GAN Diffusion, GAN vs. Diffusion, GAN
@@
To support our goal of studying generalization in generated image detection, we collect a dataset of images sampled from a wide range of models (Fig. 2). Our dataset consists of: (a) a large and systematically collected set of “in-the-wild” text-to-image latent diffusion models obtained from a model-sharing website, (b) hand-selected models from other open source architectures, and (c) closed and open state-of-the-art commercial models. We also pair these generated images with real images from other datasets. For all images in our dataset, we preserve the original image format whenever possible, without any additional compression or resizing. This is to mitigate potential bias and performance degradation in out-of-domain settings due to unwanted artifacts [43]. Our dataset contains significantly more models than previous works (Tab. 1) and spans a wider range of architectures, processing pipelines, and semantic contents.
@@ 3.1 Systematically collecting generative models
We perform our systematic collection using publicly available, open source [^2] We use the term “open source” to refer to models with public weights and source code, even if the models may be closed in some respects (e.g., private training data). models that use the Hugging Face diffusers library [124, 37] because: 1) it is a popular library for creating text-to-image models and is widely used by hobbyists, 2) thousands of such models are publicly indexed, and 3) it provides a standard interface by which we can sample images. We process them in the order of popularity, as indicated by the number of downloads. Our pipeline downloads each model and extracts relevant hyperparameters (e.g., number of diffusion steps), sampling pipeline configurations, and metadata from the model-sharing webpage [124, 37]. We sample images using a distribution of text prompts obtained from real images (Sec. 3.3). Since experiments suggest that there are diminishing returns for sampling large numbers of images from any given model, we sample a few hundred images from each one. Images with NSFW content are filtered out using a safety checker [16]. We obtain 4763 models with approximately 403 images each from this process.
While the lack of documentation in each model and the scale of data collection make it challenging to exactly characterize the model designs in this set, they are either entirely (or almost entirely) based on latent diffusion. More specifically, we categorize models as being based on latent diffusion if they perform a denoising process on a latent representation.  [^3] We note that this definition includes latent consistency models [80, 117], which are present in our dataset. Based on this criterion and the self-reported tags, all models in our systematically collected set appear to be based on latent diffusion. While pixel-based diffusion models also use the diffusers library (e.g., DeepFloyd [25]), they were incompatible with our automated generation pipeline. We record incompatible models such as these and manually sample a portion of them to construct an out-of-distribution test set (Sec. 3.4), or as manually-chosen models used for training data (Sec. 3.2).
We show examples of sampled images in Fig. 2. In Appendix D, we provide examples of models and information from their project pages. These models generate a variety of different types of images, with various types of preprocessing. For example, a large fraction of these models adapt variations of a popular pretrained latent diffusion model, Stable Diffusion [102], to different downstream applications, and use a number of adaptation strategies (e.g., using LoRA [52]). We provide the model metadata with each image to enable other possible forensics forensics applications. We discuss these in Appendix B and provide information about image and model licenses.
@@ 3.2 Collecting images from other architectures
To ensure that our dataset contains a broader range of models, we manually select 19 models from public repositories and sample 40,738 images per model on average. We note that this number is itself on par with (or more than) prior datasets with identifiable generative models. We include several GANs (e.g., StyleGANs [62, 61, 63, 109], BigGAN [10], StyleSwin [130], GigaGAN [58], ProGAN [59], ProjectedGAN [108], GANsformer [53], SAN [118], and CIPS [7]), pixel-based diffusion models (e.g., GLIDE [89], ADM [27], and DeepFloyd [25]), latent diffusion models (e.g., VQ-Diffusion [44], Diffusion Transformers [96], and Latent Flow Matching [23]), and an autoregressive model (Taming Transformers [36])
@@ 3.3 Collecting real images
To help study how real images influence forensics models, we source real images from a variety of existing datasets: LAION [110], ImageNet [26], COCO [75], FFHQ [60], CelebA [77], MetFaces [61], AFHQ [15], Forchheim [45], IMD2020 [90], Landscapes HQ [115], and VISION [114]. [^4] Following common convention in visual forensics, we refer to these images as real images, even though they may be synthetic (e.g., containing graphic design). More precisely, our goal is to distinguish “AI-generated” versions of images from the originals.
@@ 3.5 Generating images
Unconditional models are sampled until we reach the desired number of images. For class conditional models, we sample an equal number of images per class. To sample from text-conditional models, we gather prompts from multiple sources to ensure semantic diversity. We obtain captions from real images (Sec. 3.3). We either use captions that are already present in the dataset (when available), or we use BLIP [70] to generate them. The captions are then used to sample synthetic images. Some models such as GigaGAN [58] and HDiT [19] do not provide a pretrained model, so we instead use their pre-generated images. Generated images are saved in PNG format to avoid compression artifacts. However, Firefly [4] generated images are saved in JPEG format as their web UI does not allow downloading in PNG.
@@ 4 Experiments
Model Evaluation Set (mAP) Evaluation Set (Acc) Wang et al. [126] Ojha et al. [91] SB [9] GenImage [133] Ours Wang et al. [126] Ojha et al. [91] SB [9] GenImage [133] Ours Comp. Public Comp. Public Wang et al. [126] 0.897 0.696 0.516 0.642 0.535 0.600 0.714 0.527 0.508 0.533 0.509 0.517 Ojha et al. [91] 0.939 0.957 0.620 0.797 0.630 0.656 0.791 0.821 0.532 0.641 0.543 0.548 GenImage [133] 0.929 0.984 0.813 0.999 0.938 0.968 0.795 0.966 0.719 0.990 0.857 0.886 Ours 0.964 0.991 0.904 0.990 0.979 0.977 0.873 0.950 0.818 0.946 0.895 0.888 Ours - High res. 0.967 0.996 0.974 0.998 0.991 0.994 0.901 0.970 0.908 0.957 0.925 0.912
Table 2: Generalization of AI-generated image detectors across datasets. We evaluate the classifiers trained on our dataset on several datasets, including our own. We also evaluate several previously released classifiers. Our Comprehensive set (abbreviated as Comp.) pairs the generated images with original real data; the Public set pairs them with openly licensed COCO [75] and FFHQ [60] for license-compliant redistribution of the evaluation set (Sec. 3.4). We use plain CLIP-ViT-S [30, 55, 100] architecture with $224^{2}$ and $384^{2}$ (High res.) input resolutions, Wang et al. [126] and GenImage [133] use ResNet-50 [46] with $224^{2}$ input resolution, and Ojha et al. [91] uses CLIP-ViT-L with $224^{2}$ input resolution as the backbone. Our classifiers show robust performance across all evaluation sets, outperforming all baselines in out-of-distribution evaluations ([126, 91, 9] and Ours) while nearly matching GenImage [133] on its in-distribution evaluation set.

### v2

@@ Abstract
One of the key challenges of detecting AI-generated images is spotting images that have been created by previously unseen generative models. We argue that the limited diversity of the training data is a major obstacle to addressing this problem, and we propose a new dataset that is significantly larger and more diverse than prior works. As part of creating this dataset, we systematically download thousands of text-to-image latent diffusion models and sample images from them. We also collect images from dozens of popular open source and commercial models. The resulting dataset contains 2.7M images that have been sampled from 4803 different models. These images collectively capture a wide range of scene content, generator architectures, and image processing settings. Using this dataset, we study the generalization abilities of fake image detectors. Our experiments suggest that detection performance improves as the number of models in the training set increases, even when these models have similar architectures. We also find that increasing the diversity of the models improves detection performance, and that our trained detectors generalize better than those trained on other datasets. The dataset can be found in https://jespark.net/projects/2024/community_forensics
@@ 1 Introduction
Figure 1: Performance vs. model diversity. We use images sampled from different numbers of open source latent diffusion models in our Community Forensics dataset to train fake image detectors (shown in Fig. 2(a)). We fix the total number of images and only vary the number of models. The detector’s performance increases across all generator types as we train from more models, even though these added models are entirely latent diffusion. This improvement is largest for test images from out-of-distribution generators, such as pixel-based diffusion models or GANs.
@@
We hypothesize that the lack of diversity in training datasets is a major source of these shortcomings. Although today’s datasets often contain millions of fake images, they come from a relatively small number of generators. As a result, this data fails to capture many sources of variation that one might encounter in the wild. These limitations also make it challenging to accurately benchmark performance, since it is easy for cues that work well on one set of generators to fail on others.
To address these problems, we propose Community Forensics, a dataset that is significantly more diverse and comprehensive than those in prior works (Fig. 2). Our dataset contains images generated by: (a) thousands of systematically downloaded open-source latent diffusion models, (b) hand-selected open source models with various architectures, and (c) state-of-the-art commercial models.
To acquire large numbers of models, we sample images from thousands of text-to-image diffusion models hosted on a popular model-sharing website, Hugging Face [34]. We exploit the fact that these models use a common programming library [127] and thus can be sampled in a standardized way. A large fraction of them are extensions of Stable Diffusion [104], but collectively capture a variety of common model variations, such as in the architecture, image processing, and image content. We also sample images from many other open source models, including GANs [37], autoregressive models [39], and consistency models [118, 80]. To ensure sufficient diversity in the image content of the generated images, we sample generators using captions sourced from a mix of several existing datasets whenever possible.
@@
We use this dataset to study generalization in the generated image detection problem. Our experiments confirm that increasing the diversity of data improves generalization, consistent with previous findings in other areas of computer vision [111, 112, 48]. In contrast to other work, our dataset aims to increase diversity via number of generators, an underexplored axis of diversity. Through experiments, we find:
@@
  • Adding more generative models improves generatlization. Fig. 1 demonstrates the performance of fake image detection when trained on samples from varying numbers of diffusion models. Notably, the performance improves as more models are added, even across different architectures.
  • Including diverse generative model architectures significantly improves results, since classifiers do not fully generalize between generator architectures.
@@ 2 Related work
A number of datasets have been proposed for specifically detecting “deepfake” images containing manipulated faces [105, 106, 25, 67, 71, 63, 138]. Rather than focusing on face manipulation, we address creating general-purpose methods that can detect images that have been directly produced by generative models. Wang et al. [129] proposed a widely-used dataset of CNN-generated images, mixing images from GANs [58, 59, 8, 136, 12, 94] with other models [106, 10, 70, 18, 9]. This work showed that forensics models generalize between generative models, providing motivation for training on large datasets of diverse generators. However, their classifier was trained on images from a single GAN and was highly sensitive to data augmentation parameters, and more recent work shows that it does not generalize to newer models [14, 90]. Ojha et al. [90] introduced a dataset of recent diffusion models and found that training a linear classifier on CLIP features [100] extracted from ProGAN-generated images performed well. Cozzolino et al. [15] extend this work by studying the performance of CLIP-based detectors on various generative models and datasets. Epstein et al. [32] simulated detecting fake images in an online way by training a detector up to a certain year and testing it on generators released after that year. Zhu et al. [137] collected 1.4M generated images from 8 different generators. These datasets, however, only consider a handful of models (less than 20 each), limiting the generalization of their detectors. Asnani et al. proposed RED116 [5], collecting 116K images from 116 generative models for predicting the model hyperparameters of a given generated image. Guo et al. [43] extends RED116 to 140 generative models with 140K images total. We improve upon these works by collecting much more diverse generative models to improve the performance and generalization of the detector. In concurrent work, Hong et al. [49] acquires user-created images from Midjourney and CivitAI. This strategy is complementary to ours: while it aims to collect in-the-wild fake images, its distribution is centered on images that users share, and the models are not necessarily identifiable, making it challenging to rigorously analyze the dataset’s contents and to interpret experiments conducted on it.
@@
Classic work on image forensics relied on methods based on image statistics [99] and physical constraints [56], rather than learning. A number of datasets have been created for detecting images that have been manipulated using traditional methods, such as with photo editors [26, 53, 21, 64, 87]. Recent works focus on detecting synthetic images by inspecting the generator fingerprints. Zhang et al. [135] and Marra et al. [81] proposed identifying the spatial fingerprints left by the generator to detect synthetic images. Others focus on spectral anomalies to detect synthetic images. Durall et al. [29] and Dzanic et al. [30] identified that CNN-generated images fail to reproduce certain spectral properties of real images. Corvi et al. [14] study the frequency fingerprints of the generated images and analyzes the cross-architecture generalization of the detector. Bammey [7] uses high-frequency artifacts to detect generated images. However, these approaches may be brittle since the artifacts they rely on can be eliminated by post-processing [15]. We instead approach this problem in a data-driven manner, scaling the number of models, images, and architectures. Recent work has created ensembles of fake image classifiers [50]. In parallel, researchers have detected text generated by language models using supervised learning and heuristics [85, 55, 6, 125, 35, 107, 65, 117], which closely resemble those in visual forensics. However, no existing techniques that we are aware of aim to collect comprehensive datasets of community-created generators.
@@
Our work is related to the out-of-distribution recognition problem as it involves generalizing to unseen generators and image processing pipelines. A variety of approaches have been proposed for this problem, based on likelihood ratios [102, 72, 132], self-supervision [113, 86, 47, 128], internal model statistics [46, 108], temperature scaling [3, 73], and via energy-based models [28, 75, 31]. Works by Schuhmann et al. [112] and Hendrycks et al. [48] show that diverse training data and data augmentation are important to improving the robustness to out-of-distribution samples. Our results are in line with these conclusions, as we find that a diverse set of generative models and stronger augmentations improve generalization.
@@ 3 The Community Forensics Dataset
GenImage [137] 8 1.4M Diffusion, GAN Diffusion, GAN vs. ImageNet [23]
RED116 [5] 116 116K GAN, VAE, Autoregressive, … Many vs. Many
RED140 [43] 140 140K Diffusion, GAN, VAE, … Many vs. Many
@@
To support our goal of studying generalization in generated image detection, we collect a dataset of images sampled from a wide range of models (Fig. 2). Our dataset consists of: (a) a large and systematically collected set of “in-the-wild” text-to-image latent diffusion models obtained from a model-sharing website, (b) hand-selected models from other open source architectures, and (c) closed and open state-of-the-art commercial models. We also pair these generated images with real images from other datasets. We preserve the original image format where possible, without any additional compression or resizing. This is to mitigate potential bias and performance degradation in out-of-distribution settings due to unwanted artifacts [40, 42, 101]. Our dataset contains significantly more models than previous works (Tab. 1) and spans a wider range of architectures, processing pipelines, and semantic content.
@@ 3.1 Systematically collecting generative models
We perform our systematic collection using publicly available, open source [^1] We use “open source” to describe models with public weights and code, even if they may be closed in some aspect (e.g., private training data). models that use the Hugging Face diffusers library [127, 34] because: 1) it is a popular library for creating text-to-image models and is widely used by hobbyists, 2) thousands of such models are publicly indexed, and 3) it provides a standard interface by which we can sample images. We process the models in the order of popularity, as indicated by the number of downloads. Our pipeline downloads each model and extracts relevant hyperparameters (e.g., number of diffusion steps and guidance scale), sampling pipeline configurations, and metadata, from both the model cards on the model-sharing webpage and the metadata retrieved by the diffusers library [127, 34]. We sample images using a distribution of text prompts obtained from various real datasets (Sec. 3.3). Since experiments suggest diminishing returns for repeatedly sampling from any given model, we sample a few hundred images from each one. We sample 4763 models with approximately 403 images each, for a total of 1.9M images from this process.
While the lack of documentation in each model and the scale of data collection make it challenging to exactly characterize the model designs in this set, they appear to be entirely (or almost entirely) based on latent diffusion. More specifically, we categorize models as being based on latent diffusion if they perform a denoising process on a latent representation. [^2] We note that this definition includes latent consistency models [79]. Based on this criterion and the self-reported tags, all models in our systematically collected set appear to be based on latent diffusion. While pixel-based diffusion models also use the diffusers library (e.g., DeepFloyd [22]), they were incompatible with our automated generation pipeline. We record such incompatible models and manually sample them to either construct an out-of-distribution test set (Sec. 3.4), or as manually-chosen models for training data (Sec. 3.2).
We show examples of the sampled images in Fig. 2. In Appendix D, we provide examples of models and information from their project pages. These models generate a variety of images, with various types of semantic content and preprocessing. For example, a large fraction of these models adapt variations of a popular pretrained latent diffusion model, Stable Diffusion [104], to different downstream applications, and use a number of adaptation strategies (e.g., using LoRA [51]). We provide the model metadata with each image to enable other possible forensics applications. We discuss these in Appendix B and provide information about image and model licenses.
@@ 3.2 Collecting images from other architectures
To ensure that our dataset contains a broader range of models, we manually select 19 models from public repositories and sample an average of 40,738 images per model, resulting in 774K images total. We note that this number is itself on par with (or more than) many prior datasets with identifiable generative models [129, 90, 32, 15, 7, 137]. We include several GANs (e.g., StyleGANs [61, 60, 62, 110], BigGAN [8], StyleSwin [134], GigaGAN [57], ProGAN [58], ProjectedGAN [109], GANsformer [52], SAN [119], and CIPS [4]), pixel-based diffusion models (e.g., GLIDE [88], ADM [24], and DeepFloyd [22]), latent diffusion models (e.g., VQ-Diffusion [41], Diffusion Transformers [96], and Latent Flow Matching [20]), and an autoregressive model (Taming Transformers [33]).
@@ 3.3 Collecting real images
To help study how real images influence forensics models, we source real images from a variety of existing datasets: LAION [111], ImageNet [23], COCO [74], FFHQ [59], CelebA [76], MetFaces [60], AFHQ [13], Forchheim [44], IMD2020 [89], Landscapes HQ [116], and VISION [115]. [^3] Following common convention, we refer to these images as real images, even though they may be synthetic (e.g., graphic design). More precisely, our goal is to distinguish “AI-generated” images from the originals.
@@ 3.5 Generating images
Unconditional models are sampled until we reach the desired number of images. For class conditional models, we sample an equal number of images per class. Text-conditional models are sampled using captions obtained from real images (Sec. 3.3). We either use captions that are already present in the dataset or use BLIP [69] to generate them. Some models such as GigaGAN [57] and HDiT [16] do not provide a pretrained model, so we instead use their pre-generated images. Generated images are saved in PNG format to avoid compression artifacts. However, Firefly [1] generated images are saved in JPEG format as their web UI only allows downloading in JPEG.
@@ 4 Experiments
[…]
Table 2: Generalization of AI-generated image detectors across benchmarks. We evaluate the classifiers trained on our dataset on several benchmarks, including our own. We also evaluate several previously released classifiers. Our Comprehensive set (abbreviated as Comp.) pairs the generated images with original real data; the Public set pairs them with openly licensed COCO [74] and FFHQ [59] for license-compliant redistribution of the evaluation set (Sec. 3.4). We use plain CLIP-ViT-S [27, 54, 100] architecture with $224^{2}$ and $384^{2}$ (High res.) input resolutions, Wang et al. [129] and GenImage [137] use ResNet-50 [45] with $224^{2}$ input resolution, and Ojha et al. [90] uses CLIP-ViT-L with $224^{2}$ input resolution as the backbone. Our classifiers show robust performance across all evaluation sets, outperforming all baselines in out-of-distribution settings ([129, 90, 7] and Ours) and nearly matches GenImage [137] on its in-distribution evaluation set. $\blacklozenge$: RED140 [43] is trained following our training procedure with $384^{2}$ input resolution as they do not provide a trained classifier.